(15 points) Prove by induction that for an odd k > 1, the number 2n+2 divides k2 – 1 for all every positive integer n.

induction question, thanks.
Answer #1

so K² - 1 = 2 M - for some MEN. 0 for Now all • we will prove it is true n +-1 h) 2 K²_1 = K² -1 = K - (k² + 1) (K²-1) 1 - b